# Aligning Perception, Reasoning, Modeling and Interaction: A Survey on Physical AI

### v1

@@ A Introduction
A particularly promising emerging pathway leverages large language models (LLMs) and multimodal large language models (MLLMs). Recent breakthroughs have demonstrated their impressive capability in learning from massive datasets and performing sophisticated end-to-end reasoning across multiple modalities [28, 29, 30, 31, 32, 33]. Notably, works employing multi-agent systems and tool integration have achieved performance comparable to top human competitors in physics competitions [34, 35], suggesting the potential for more general physical understanding.
@@
Contributions. We present a comprehensive three-tier taxonomy that systematically captures reasoning capabilities and downstream applications in the current landscape of physical AI systems. We categorize contemporary research efforts into four fundamental capabilities: physical perception (drawing conclusions from observations of physical phenomena), physics reasoning (utilizing symbolic representations and mathematical methods to solve theoretical physics problems), world modeling (modeling and predicting dynamics in virtual environments that mirror our physical world), and embodied interaction (navigating and manipulating real objects in physical environments).
@@ D.1 Benchmarking Physics Reasoning
Figure 4: Multi-agent physics reasoning system (Figure used courtesy of [34]).
@@ D.2 General Physics Reasoning Models
A significant leap is the development of autonomous agentic systems that manage entire scientific workflows. A key example is The AI Scientist, which generates complete research papers from ideation to final manuscript at minimal cost [141]. Its successor, The AI Scientist-V2, was the first AI system to have a paper accepted through formal peer review [142]. Other agents demonstrate performance equal to human experts in high-level competitions. A major breakthrough came with Physics Supernova (Figure 4), an AI agent that solved physics problems at the level of top human students. It ranked 14th out of 406 contestants at the 2025 International Physics Olympiad with a score of 23.5 out of 30 [34].
Multi-agent systems have also shown great success in specific scientific domains. In these systems, several specialized AI agents collaborate. For instance, AtomAgents helps design novel metal alloys [143]. The Multimodal Universe project processes vast astronomical datasets and has compiled over 100 TB of data for machine learning research [144]. Visual and multimodal reasoning remains a critical challenge. One successful approach used text descriptions of images to help AI better understand physics diagrams. This system won first place at the ICML 2025 AI4Math Challenge [35]. These advancements show a clear trajectory toward fully automated science labs. However, the field still relies on specialized systems, particularly for tasks with strong visual dependence. A single foundational model for general physical reasoning does not yet exist.
@@ D.3 Theoretical Physics Solvers
Data-driven approaches aim to discover unknown physical laws directly from data. A prominent example is Symbolic Regression (SR), which extracts interpretable mathematical expressions from observations without prior knowledge of the governing equations. This makes SR a powerful tool for generating human-readable models that overcome the "black-box" nature of many neural networks. The primary obstacle for SR is navigating its vast combinatorial search space while ensuring physical plausibility. To this end, recent advancements focus on incorporating hard physical constraints, such as unit consistency, to accelerate the learning process and guide the discovery toward physically meaningful equations [146].
@@ E.1 Generative Models
Modern video generation models, such as diffusion-based approaches like GPT4Motion [164] and VLIPP [174], are progressively constrained by physics laws, making the generated video sequences not only visually coherent but also physically plausible. For example, as illustrated in Figure 5, GPT4Motion is constrained by physics laws through physical simulators like PyBullet [175] by using GPT-4 to generate Blender scripts that simulate realistic physics dynamics, ensuring both visual consistency and adherence to physical laws across video frames, while VLIPP employs chain-of-thought reasoning in vision-language models to predict physically plausible motion trajectories that guide diffusion models to generate temporally coherent and physically accurate video sequences.

### v2

@@ Abstract
“What I cannot create, I do not understand.”
Richard Feynman [^1] Richard Feynman (1918-1988) was a Nobel Prize-winning physicist known for his groundbreaking work in quantum electrodynamics and the invention of Feynman diagrams.
@@ A Introduction
Figure 1: Overview of four physical understanding capabilities of current AI systems. Physical Perception denotes model’s ability to recognize the properties of objects; Physics Reasoning refers to model’s capability to understand symbolic systems and physical laws; World Modeling represents its capacity for simulation and prediction; and Embodied Interaction characterizes model’s ability to interact with the physical environment.
@@
A particularly promising emerging pathway leverages large language models (LLMs) and multimodal large language models (MLLMs). Recent breakthroughs have demonstrated their impressive capability in learning from massive datasets and performing sophisticated end-to-end reasoning across multiple modalities [28, 29, 30, 31, 32, 33]. Notably, works employing multi-agent systems and tool integration have achieved performance comparable to top human competitors in physics competitions [qiu2024physics, 34], suggesting the potential for more general physical understanding.
@@
Contributions. We present a comprehensive three-tier taxonomy that systematically captures reasoning capabilities and downstream applications in the current landscape of physical AI systems. We categorize contemporary research efforts into four fundamental capabilities which are shown in Figure 1: physical perception (drawing conclusions from observations of physical phenomena), physics reasoning (utilizing symbolic representations and mathematical methods to solve theoretical physics problems), world modeling (modeling and predicting dynamics in virtual environments that mirror our physical world), and embodied interaction (navigating and manipulating real objects in physical environments).
@@ D.1 Benchmarking Physics Reasoning
Figure 5: Multi-agent physics reasoning system across 3 use case spanning theoretical, computational and experimental research.
@@ D.2 General Physics Reasoning Models
A significant leap is the development of agentic systems. A key example is The AI Scientist, which generates research papers from ideation to final manuscript [140]. Its successor, The AI Scientist-V2, was the first AI system to have a paper accepted through peer review [141]. Other agents demonstrate performance equal to human experts in high-level competitions. In particular, Physics Supernova [142] ranked 14th out of 406 contestants at the 2025 International Physics Olympiad (IPhO).
Multi-agent systems have also shown great value for scientific discovery [143] as we demonstrate in Figure 5. For instance, AtomAgents helps design novel metal alloys [144]. The Multimodal Universe project processes vast astronomical datasets and has compiled over 100 TB of data for machine learning research [145]. Visual and multimodal reasoning remains a critical challenge. One successful approach used text descriptions of images to help AI better understand physics diagrams, which won first place at the ICML 2025 AI4Math Challenge [34]. These advancements show a clear trajectory toward automating scientific workflows. However, the field still relies on specialized systems, particularly for tasks with strong visual dependence.
@@ D.3 Theoretical Physics Solvers
Data-driven approaches aim to discover unknown physical laws directly from data. A prominent example is Symbolic Regression (SR), which extracts interpretable mathematical expressions without prior knowledge of the governing equations. This makes SR a powerful tool for generating human-readable models that overcome the "black-box" nature of many neural networks. The primary obstacle for SR is navigating its vast combinatorial search space while ensuring physical plausibility. To this end, recent advancements focus on incorporating hard physical constraints, such as unit consistency, to accelerate the learning process and guide the discovery toward physically meaningful equations [147].
@@ E.1 Generative Models
Modern video generation models, such as diffusion-based approaches like GPT4Motion [165] and VLIPP [175], are progressively constrained by physics laws, making the generated video sequences not only visually coherent but also physically plausible. For example, as illustrated in Figure 6, GPT4Motion is constrained by physics laws through physical simulators like PyBullet [176] by using GPT-4 to generate Blender scripts that simulate realistic physics dynamics, ensuring both visual consistency and adherence to physical laws across video frames, while VLIPP employs chain-of-thought reasoning in vision-language models to predict physically plausible motion trajectories that guide diffusion models to generate temporally coherent and physically accurate video sequences. DINO-Foresight [177] supports multiple future understanding tasks and avoids reconstructing irrelevant details of pixel based methods.